derive an equation for the kinetic energy of a cart going down an inclined ramp in terms of h, l, g and M. the answer is KE= E-m(h/l)(g) but how do you get that answer?? thanks in advance
Answer #1

using conservation of energy

1/2 m 02 + m g h = 1/2 m v2 + m g x 0

v2 = 2 g h

h = L cos

h = ( L2 - B2 )1/2

KE = 1/2 m v2

= 1/2 m 2 g (L2 - B2)1/2

= m g (L2 - B2)1/2

so KE = m g L cos

or

KE = m g (L2 - B2)1/2